# Linear Attention for Joint Power Optimization and User-Centric Clustering in Cell-Free Networks

### v1

@@ III-A Overview of CosFormer
The standard Transformer [vaswani2017attention] relies on an attention mechanism that captures dependencies among tokens (UEs in our case), but its quadratic complexity limits scalability. CosFormer [qin2022cosformer] overcomes this issue by introducing a linearized attention mechanism that preserves accuracy while significantly reducing computational cost. Among linear Transformer variants [choromanski2020rethinking, wang2020linformer, beltagy2020longformer], CosFormer is particularly well suited to our setting due to the following properties:
@@
    Computational efficiency: Through asymmetric projections, CosFormer attains linear complexity, reducing memory use and accelerating training, an essential feature for scalability in large cell-free networks.
Building on this foundation, we introduce our modified architecture, ELU-CosFormer, which augments CosFormer [qin2022cosformer] with pilot-aware constraints and refined activation functions. These enhancements enable the model to jointly predict AP clusters and powers while eliminating pilot contamination and ensuring SE fairness. Once trained, ELU-CosFormer produces clusters and UL/DL power levels directly from the spatial positions of the UEs and APs, allowing seamless adaptation to varying network configurations without the need for retraining.
@@
In this section, we introduce the proposed learning model, ELU-CosFormer, which performs joint AP clustering and UL/DL power allocation. We first describe the overall model architecture and subsequently detail the linearized attention mechanism based on the modified CosFormer formulation.
@@ IV-A Model architecture
    with trainable parameters $\mathbf{W}_{\text{in}}$ and $\mathbf{b}_{\text{in}}$. Through this dynamic embedding, the model handles seamlessly varying network sizes without architectural changes.
@@
    Encoder layers: The embedded tensor $\mathbf{Z}^{(0)}=\mathbf{Z}_{\text{in}}\in\mathbb{R}^{K\times d_{\text{mod}}}$ is processed through $M$ encoder layers. Each layer consists of a modified CosFormer multi-head attention (MHA), denoted by $\mathcal{A}(\cdot)$, followed by a feed-forward network (FFN). Moreover, the encoder layers apply layer normalization and residual connections to stabilize training and preserve information across layers [qin2022cosformer]. Specifically, the $m$-th encoder layer applies the following operations (see Figure 3):
@@
    $\mathbf{H}^{(m)}$ $=\text{ReLU}\!\left(\tilde{\mathbf{Z}}_{1}^{(m)}\mathbf{W}_{1}^{(m)\top}+\mathbf{b}_{1}^{(m)}\right),$ (13)
    $\mathbf{Z}^{(m)}$ $=\mathbf{Z}_{1}^{(m)}+\mathbf{H}^{(m)}\mathbf{W}_{2}^{(m)\top}+\mathbf{b}_{2}^{(m)},$ (14)
    where $\tilde{\mathbf{Z}}$ denotes a layer-normalized version of a tensor $\mathbf{Z}$; $\mathbf{W}_{1}^{(m)}$, $\mathbf{b}_{1}^{(m)}$, $\mathbf{W}_{2}^{(m)}$, and $\mathbf{b}_{2}^{(m)}$ are trainable parameters; ReLU refers to an activation function [rasamoelina2020review] introducing a non-linearity.
    All $M$ layers share the same structure and the output of each encoder layer preserves the dimension $\mathbb{R}^{K\times d_{\text{mod}}}$, serving as input to the next layer. After $M$ layers, the final tensor $\mathbf{Z}^{(M)}=\mathbf{Z}_{\text{out}}$ represents the encoder output that will be processed next by the output heads. The attention mechanism in (12) is detailed in Sect. IV-B.
@@
    • User-centric clustering mask: A fully-connected layer with sigmoid activation [rasamoelina2020review] produces an initial mask tensor $\mathbf{M}\in\mathbb{R}^{K\times L}$, where each entry represents the association strength between a UE and an AP:
      $\mathbf{M}=\text{sigmoid}\left(\mathbf{Z}_{\text{out}}\mathbf{W}_{\text{mask}}+\mathbf{b}_{\text{mask}}\right),$ (15)
@@
      To mitigate pilot contamination, a two-step procedure is applied. First, for each AP, only the top $\tau_{p}$ UEs with the highest mask scores are retained via $\text{Top}_{\tau_{p}}(\mathbf{M})$. Then, a thresholding operation assigns the AP to serve the UE if the retained score exceeds a threshold $\xi$, yielding the final binary clustering mask:
      $\mathbf{M}_{\tau_{p}}=\mathbbm{1}\left(\text{Top}_{\tau_{p}}(\mathbf{M})\odot\mathbf{M}>\xi\right).$ (16)
@@
      $\tilde{p}_{k}^{\text{UL}}=\text{sigmoid}\left(\mathbf{Z}_{\text{out}}\mathbf{W}_{\text{UL}}+\mathbf{b}_{\text{UL}}\right),$ (17)
    • Downlink power prediction: Another fully-connected layer with sigmoid activation predicts normalized DL powers:
      $\tilde{p}_{k}^{\text{DL}}=\text{sigmoid}\left(\mathbf{Z}_{\text{out}}\mathbf{W}_{\text{DL}}+\mathbf{b}_{\text{DL}}\right),$ (18)
      where $\mathbf{W}_{\text{UL}},\mathbf{W}_{\text{DL}},\mathbf{b}_{\text{UL}},\mathbf{b}_{\text{DL}}$ are trainable parameters.
    The predicted powers are then denormalized and rescaled to produce the final output $\widehat{\mathbf{p}}\in\mathbb{R}^{K\times 2}$:
@@
    $\hat{p}_{k}^{\text{DL}}$ $=\frac{\check{p}_{k}^{\text{DL}}\sum_{l=1}^{L}\overline{P}_{l}^{\text{DL}}}{\sum_{k=1}^{K}\check{p}_{k}^{\text{DL}}},$ (20)
    with
    $\check{p}_{k}^{\text{DL}}$ $=\Delta_{\text{DL}}\tilde{p}_{k}^{\text{DL}}+\underline{P}_{\text{DL}},$ (21)
    where $\Delta_{\text{UL}}=\overline{P}_{\text{UL}}-\underline{P}_{\text{UL}}$ and $\Delta_{\text{DL}}=\overline{P}_{\text{DL}}-\underline{P}_{\text{DL}}$ denote the UL and DL power ranges, respectively. Here, $\overline{P}_{\text{UL}}$ and $\underline{P}_{\text{UL}}$ are the maximum and minimum UL powers for UE $k$, while $\overline{P}_{\text{DL}}$ and $\underline{P}_{\text{DL}}$ are the corresponding DL bounds. The DL rescaling ensures that the total DL power budget across all APs is preserved.
@@ IV-B Attention mechanism
We now detail the attention mechanism used in (12). The queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) [vaswani2017attention] are simultaneously computed from the layer-normalized input $\tilde{\mathbf{Z}}_{\text{in}}$ using a single linear transformation:
$\text{QKV}=\tilde{\mathbf{Z}}_{\text{in}}\mathbf{W}_{\text{qkv}}+\mathbf{b}_{\text{qkv}},$ (22)
where $\mathbf{W}_{\text{qkv}}\in\mathbb{R}^{d_{\text{mod}}\times 3d_{\text{mod}}}$ and $\mathbf{b}_{\text{qkv}}\in\mathbb{R}^{3d_{\text{mod}}}$ are trainable parameters. The resulting tensor $\text{QKV}\in\mathbb{R}^{B\times K\times 3d_{\text{mod}}}$ is split along the last dimension into three parts, yielding $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times K\times d_{\text{mod}}}$. These are then reshaped into $N_{h}$ attention heads of dimension $d_{\text{head}}=d_{\text{mod}}/N_{h}$, giving
@@
with $\phi(\cdot)$ being an activation function chosen to approximate the softmax kernel [rasamoelina2020review] of the original Transformer while enabling linearized attention.
@@
where $\mathbf{W}\in\mathbb{R}^{K\times K}$ is a fixed cosine reweighting matrix with entries $w_{ij}=\cos\!\left(\frac{\pi(i-j)}{2K}\right)$. This modulation biases attention toward nearby tokens, mimicking the locality bias of softmax without computing full pairwise interactions, thereby avoiding the quadratic complexity. In our context, this allows the model to distinguish between nearby and distant UEs/APs without introducing additional trainable parameters.
@@
$\mathcal{A}(\tilde{\mathbf{Z}}_{\text{in}})=\mathbf{A}\mathbf{W}_{\text{out}}^{\top}+\mathbf{b}_{\text{out}},$ (27)
@@
Our choice of activation function. In the original CosFormer [qin2022cosformer], the kernel feature mapping function $\phi(\cdot)$ is implemented using ReLU. In this work, we propose a simple yet efficient modification: replacing ReLU with the ELU [rasamoelina2020review]. This choice is motivated by both physical insights and empirical evidence. Formally, the two activations are defined as [rasamoelina2020review]:
$\begin{aligned} \text{ReLU}(x)&=\max(0,x),\\ \text{ELU}(x)&=\begin{cases}x&\textrm{if $x>0$},\\ \alpha(e^{x}-1)&\textrm{if $x\leq 0$},\end{cases}\end{aligned}$ (28)
@@
ELU provides smoother gradients and nonzero outputs for negative inputs, as illustrated in Figure 4, thereby preserving richer feature interactions and improving gradient flow and numerical stability during training. This is particularly important in our setting, where spatial coordinates require the attention mechanism to capture subtle UE-AP variations. In contrast, ReLU zeroes out all negative inputs, potentially discarding useful spatial information and leading to sparse gradients that hinder convergence. By deliberately allowing negative values, ELU maintains continuity across the input domain and prevents inactive attention heads, enabling the attention mechanism to exploit weak but meaningful AP-UE associations that would otherwise be suppressed. Such associations are critical under the max-min fairness criterion, where low‑gain links can influence the performance of the worst‑case user.
Moreover, the exponential tail of ELU produces a differentiable kernel that interacts harmoniously with the cosine modulation in (26), resulting in smoother and more stable attention distributions. This behavior aligns with the naturally smooth spatial variations in wireless propagation, leading to physically realistic clustering boundaries and balanced power allocation.
Figure 4: Comparison of ReLU and ELU activation functions. ELU smoothly transitions for negative inputs using an exponential curve, improving gradient flow and stability.
To validate this design choice, we evaluate the per-UE SE achieved by a CosFormer using several activations [rasamoelina2020review] with the simulation parameters in Table III. The results reported in Figures 5 demonstrate that ELU consistently yields higher SE across all user counts $K$ for both UL and DL scenarios, confirming that ELU provides a more effective kernel transformation for our problem.
(a) UL
(b) DL
@@ V-A Dataset construction
To train and evaluate the learning model, we construct a synthetic dataset that captures diverse cell‑free network configurations and realistic deployment conditions. The dataset construction follows four main steps:
@@ 1. Step 1)
    • random 2D positions within a bounded area for $K$ single‑antenna UEs, denoted as $\{\mathbf{u}_{k}\}_{k=1}^{K}$, with $\mathbf{u}_{k}\in\mathbb{R}^{2}$, and $L$ uniformly distributed APs, denoted as $\{\mathbf{a}_{l}\}_{l=1}^{L}$, with $\mathbf{a}_{l}\in\mathbb{R}^{2}$;
    • Large‑scale fading coefficients $\beta_{lk}$ computed as in [miretti2022closed], which are required later by the closed‑form solution to solve the max-min SE optimization problems in (9) and (10) for each set of clusters.
@@ 2. Step 2)
    Noise injection: To simulate localization errors and improve robustness, Gaussian noise is added to the UE and AP positions:
@@
    $\mathbf{a}^{\prime}_{l}$ $=\mathbf{a}_{l}+\delta_{l},\quad\delta_{l}\sim\mathcal{N}(0,\sigma_{e}^{2}),$ (30)
    where $\sigma_{e}=1\,\mathrm{m}$, following [xue2021analysis]. This perturbation accounts for position estimation uncertainty in practice.
@@ 3. Step 3)
    Normalization: All features are normalized to the range $[0,1]$ using min-max scaling. For a feature vector $\boldsymbol{\xi}$:
    $\boldsymbol{\tilde{\xi}}=\frac{\boldsymbol{\xi}-\min(\boldsymbol{\xi})}{\max(\boldsymbol{\xi})-\min(\boldsymbol{\xi})+\varepsilon},$ (31)
@@
    The normalized noisy positions are structured into an input ${\bf X}\in\mathbb{R}^{K\times(2L+2)}$, which feeds the model a sequence of $K$ tokens. Each token represents one UE together with its relative information to all APs, aligning naturally with the token‑based processing paradigm of Transformers. This design allows the attention mechanism to capture dependencies across users and identify which APs are most relevant for serving each UE, while ensuring generalization across different network sizes and topologies.
@@
    1:Training dataset $\{\mathbf{X}_{i}\}_{i=1}^{S}$, batch size $S$, trade‑off parameter $\lambda$
@@
    6:   Compute powers by solving (9) and (10) [miretti2022closed]

### v2

@@ III-A Overview of CosFormer
The standard Transformer [35] relies on an attention mechanism that captures dependencies among tokens (UEs in our case), but its quadratic complexity limits scalability. CosFormer [32] overcomes this issue by introducing a linearized attention mechanism that preserves accuracy while significantly reducing computational cost. Among linear Transformer variants [12, 36, 4], CosFormer is particularly well suited to our setting due to the following properties:
@@
    Computational efficiency: CosFormer attains linear complexity, reducing memory use and accelerating training, an essential feature for scalability in large CF networks.
Building on this foundation, we introduce our modified architecture, ELU-CosFormer, which augments CosFormer [32] with pilot-aware constraints and modified activation functions. These enhancements enable the model to jointly predict AP clusters and powers while eliminating pilot contamination and ensuring SE fairness. Once trained with data that include channel modeling and estimation, ELU-CosFormer produces clusters and UL/DL power levels directly from the spatial positions of the UEs and APs, allowing seamless adaptation to varying user loads.
@@
In this section, we introduce the proposed learning model, ELU-CosFormer, which performs joint AP clustering and UL/DL power allocation. We first describe the overall model architecture, and subsequently detail the linearized attention mechanism based on the modified CosFormer formulation.
@@ IV-A Model architecture
    with trainable parameters $\mathbf{W}_{\text{in}}$ and $\mathbf{b}_{\text{in}}$. The input is a sequence of $K$ tokens, where each token corresponds to one UE. Since the same projection is applied independently to every token, the operation does not assume a fixed number of UEs. Consequently, changes in $K$ only modify the sequence length, while the model parameters remain unchanged. The subsequent encoder processes this sequence using shared weights among UEs and preserves the sequence token-structure, enabling the learning model to handle varying user loads seamlessly without requiring architectural modifications.
@@
    Encoder layers: The embedded tensor $\mathbf{Z}^{(0)}=\mathbf{Z}_{\text{in}}\in\mathbb{R}^{K\times d_{\text{mod}}}$ is processed through $M$ encoder layers. Each layer consists of a modified CosFormer multi-head attention (MHA), denoted by $\mathcal{A}(\cdot)$, followed by a feed-forward network (FFN). Moreover, the encoder layers apply layer normalization and residual connections to stabilize training and preserve information across layers [32]. Specifically, the $m$-th encoder layer applies the following operations (see Figure 3):
@@
    $\mathbf{H}^{(m)}$ $=\mathrm{ReLU}\!\left(\tilde{\mathbf{Z}}_{1}^{(m)}\mathbf{W}_{1}^{(m)\top}+\mathbf{b}_{1}^{(m)}\right),$ (13)
    $\mathbf{Z}^{(m)}$ $=\tilde{\mathbf{Z}}_{1}^{(m)}+\mathbf{H}^{(m)}\mathbf{W}_{2}^{(m)\top}+\mathbf{b}_{2}^{(m)},$ (14)
    where $\tilde{\mathbf{Z}}$ denotes a layer-normalized version of a tensor $\mathbf{Z}$; $\mathbf{W}_{1}^{(m)}$, $\mathbf{b}_{1}^{(m)}$, $\mathbf{W}_{2}^{(m)}$, and $\mathbf{b}_{2}^{(m)}$ are trainable parameters; $\mathrm{ReLU}$ refers to an activation function [33] introducing a non-linearity.
    All $M$ layers share the same structure, and the output of each encoder layer preserves the dimension $\mathbb{R}^{K\times d_{\text{mod}}}$, serving as input to the next layer. After $M$ layers, the final tensor $\mathbf{Z}^{(M)}=\mathbf{Z}_{\text{out}}$ represents the encoder output that will be processed next by the output heads. The attention mechanism in (12)-(14) is detailed in Sect. IV-B.
@@
[…]
      $\mathbf{M}=\mathrm{sigmoid}\left(\mathbf{Z}_{\text{out}}\mathbf{W}_{\text{mask}}+\mathbf{b}_{\text{mask}}\right),$ (15)
@@
      To mitigate pilot contamination, a two-step procedure is applied. First, for each AP, only the top $\tau_{p}$ UEs with the highest mask scores are retained via $\mathrm{Top}_{\tau_{p}}(\mathbf{M})$. Then, a thresholding operation assigns the AP to serve the UE if the retained score exceeds a threshold $\xi$, yielding the final binary clustering mask:
      $\mathbf{M}_{\tau_{p}}=\mathbbm{1}\left(\mathrm{Top}_{\tau_{p}}(\mathbf{M})\odot\mathbf{M}>\xi\right).$ (16)
@@
      $\tilde{p}_{k}^{\text{UL}}=(\mathrm{sigmoid}\left(\mathbf{Z}_{\text{out}}\mathbf{W}_{\text{UL}}+\mathbf{b}_{\text{UL}}\right))_{k},$ (17)
      where $\mathbf{W}_{\text{UL}}$ and $\mathbf{b}_{\text{UL}}$ are trainable parameters. The denormalized UL powers are obtained as
@@
      where $\Delta_{\text{UL}}=\overline{P}_{\text{UL}}-\underline{P}_{\text{UL}}$ denotes the UL power range. Here, $\overline{P}_{\text{UL}}$ and $\underline{P}_{\text{UL}}$ are the maximum and minimum allowable UL transmit powers.
    • Downlink power prediction: A fully-connected layer with sigmoid activation outputs normalized per-AP DL power weights:
      $\tilde{p}_{k,l}^{\mathrm{DL}}=(\mathrm{sigmoid}\!\left(\mathbf{Z}_{\mathrm{out}}\mathbf{W}_{\mathrm{DL}}+\mathbf{b}_{\mathrm{DL}}\right))_{k,l},$ (19)
      where $\mathbf{W}_{\mathrm{DL}}$ and $\mathbf{b}_{\mathrm{DL}}$ are trainable parameters, and $\tilde{p}_{k,l}^{\mathrm{DL}}\in[0,1]$ represents the relative power weight assigned by AP $l$ to UE $k$.
    To enforce the per-AP power constraint in (10), the predicted weights are scaled by the AP power budget:
    $\hat{p}_{k,l}^{\mathrm{DL}}=\frac{\tilde{p}_{k,l}^{\mathrm{DL}}}{\sum_{i=1}^{K}\tilde{p}_{i,l}^{\mathrm{DL}}}\,\overline{P}_{l}^{\mathrm{DL}}.$ (20)
    Since the available DL power labels are defined per UE, the total predicted DL power allocated to UE $k$ is computed as $\hat{p}_{k}^{\mathrm{DL}}=\sum_{l=1}^{L}\hat{p}_{k,l}^{\mathrm{DL}}$. These values are then normalized using the min-max scaling criterion introduced in Sect. V-A (Step 3) to obtain the predicted labels used for training. This ensures that the learning model respects the per-AP power constraints while remaining consistent with the available optimal labels used in the loss function defined in Sect. V-B.
@@ IV-B Attention mechanism
We now detail the attention mechanism used in (12)-(14). The queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) [35] are simultaneously computed from the layer-normalized input $\tilde{\mathbf{Z}}_{\text{in}}$ using a single linear transformation:
$\mathbf{\Omega}=\tilde{\mathbf{Z}}_{\text{in}}\mathbf{W}_{\mathbf{\Omega}}+\mathbf{b}_{\mathbf{\Omega}},$ (21)
where $\mathbf{W}_{\mathbf{\Omega}}\in\mathbb{R}^{d_{\text{mod}}\times 3d_{\text{mod}}}$ and $\mathbf{b}_{\mathbf{\Omega}}\in\mathbb{R}^{3d_{\text{mod}}}$ are trainable parameters. The resulting tensor $\mathbf{\Omega}\in\mathbb{R}^{B\times K\times 3d_{\text{mod}}}$ is split along the last dimension into three parts, yielding $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times K\times d_{\text{mod}}}$. These are then reshaped into $N_{h}$ attention heads of dimension $d_{\text{head}}=d_{\text{mod}}/N_{h}$, giving
@@
with $\phi(\cdot)$ being an activation function chosen to approximate the softmax kernel [33] of the original Transformer while enabling linearized attention.
@@
where $\mathbf{W}\in\mathbb{R}^{K\times K}$ is a fixed cosine reweighting matrix with entries $w_{ij}=\cos\!\left(\pi(i-j)/({2K})\right)$. This modulation biases attention toward nearby tokens, mimicking the locality bias of softmax without computing full pairwise interactions, thereby avoiding the quadratic complexity. In our context, this allows the model to distinguish between nearby and distant UEs/APs without introducing additional trainable parameters.
@@
$\mathcal{A}\!\left(\tilde{\mathbf{Z}}^{(m-1)}\right)=\mathbf{A}\mathbf{W}_{\text{out}}^{\top}+\mathbf{b}_{\text{out}},$ (26)
@@
Our choice of activation function. In the original CosFormer [32], the kernel feature mapping function $\phi(\cdot)$ is implemented using rectified linear unit (ReLU). In this work, we propose a simple yet efficient modification: replacing ReLU with the ELU [33]. This choice is motivated by both physical insights and empirical evidence. Formally, the two activations are defined as [33]:
$\begin{aligned} \mathrm{ReLU}(x)&=\max(0,x),\\ \mathrm{ELU}(x)&=\begin{cases}x&\textrm{if $x>0$},\\ \alpha(e^{x}-1)&\textrm{if $x\leq 0$},\end{cases}\end{aligned}$ (27)
@@
ELU provides smoother gradients and nonzero outputs for negative inputs, preserving richer feature interactions and improving gradient flow and numerical stability during training. This is particularly important in our setting, where spatial coordinates require the attention mechanism to capture subtle UE-AP variations. In contrast, ReLU zeroes out all negative inputs, potentially discarding useful spatial information and leading to sparse gradients that hinder convergence. By deliberately allowing negative values, ELU maintains continuity across the input domain and prevents inactive attention heads, enabling the attention mechanism to exploit weak but meaningful AP-UE associations that would otherwise be suppressed. Such associations are critical under MMF criterion, where low-gain links can influence the performance of the worst-case UE.
Moreover, the exponential tail of ELU produces a smooth and fully differentiable activation, which interacts naturally with the cosine modulation in (25) (through $\bf{W}$). This combination yields a differentiable kernel that results in smoother and more stable attention distributions, reducing abrupt variations in the attention weights. This behavior is consistent with the inherently smooth spatial variations observed in wireless propagation, where signal strengths change gradually rather than abruptly. Consequently, the learned attention patterns better reflect realistic relationships between UEs and APs, leading to more physically meaningful clustering boundaries and a more balanced power allocation across the network. The effectiveness of this activation function is evaluated in Sect. VI via an ablation experiment.
@@ V-A Dataset Construction
To train and evaluate the learning model, we construct, based on the model in Sect. II, a synthetic dataset that captures diverse cell-free network configurations and realistic deployment conditions. The dataset construction follows four main steps:
@@ 1. Step 1)
    • random 2D positions within a bounded area for $K$ single-antenna UEs, denoted as $\{\mathbf{u}_{k}\}_{k=1}^{K}$, with $\mathbf{u}_{k}\in\mathbb{R}^{2}$, and $L$ uniformly distributed APs, denoted as $\{\mathbf{a}_{l}\}_{l=1}^{L}$, with $\mathbf{a}_{l}\in\mathbb{R}^{2}$;
    • large-scale fading coefficients $\beta_{lk}$, correlation matrices $\mathbf{R}_{lk}$, MMSE estimated channels $\mathbf{\widehat{h}}_{lk}$, MMSE precoders $\mathbf{w}_{k}$ and combiners $\mathbf{v}_{k}$ computed as detailed in Sect. II, which are required later by the closed-form solution to solve the max-min SE optimization problems in (9) and (10) for each set of clusters.
@@ 2. Step 2)
    Noise injection: To simulate localization errors and improve robustness, Gaussian noise is added to UE positions:
@@
    where $\sigma_{e}=1\,\mathrm{m}$, following [38]. This perturbation accounts for position estimation errors.
@@ 3. Step 3)
    Normalization: All features are normalized to the range $[0,1]$ using min-max scaling. For a feature vector $\bm{\xi}$, we use
    $\bm{\tilde{\xi}}=\frac{\bm{\xi}-\min(\bm{\xi})}{\max(\bm{\xi})-\min(\bm{\xi})+\varepsilon},$ (29)
@@
    1:Training dataset $\{\mathbf{X}_{i}\}_{i=1}^{S}$, batch size $S$, trade-off parameter $\lambda$
@@
    6:   Compute powers by solving (9) and (10) [24]